AGW Commissioning and LUCIFER Commissioning on 12-13 September 2008

Observers: J. Hill, N. Ageorges, W. Seifert, and numerous other LUCIFERians on mtn
More Observers: D. Miller and A. Rakich in Tucson
Software Support: N. Cushing on mtn
Telescope Operators: D. Gonzalez Huerta and W. Wack

Summary

The last of the clouds cleared away just after sunset. It was clear the rest of the night. Seeing was poor at first with everything out of temperature equilibrium. Our first significant problem was that when authorized as LUCIFER, we don't find off-axis guide stars. A guide star 90 sec due south of the target lands at the very upper right corner off the acquisition image (an error of ~15 arcsec in 90 sec). The corresponding error when authorized at IRTC is a factor of 4-5 smaller. ??? The only differences in the PCSInstrument.conf file is that IRTC has wavelength 1.4 versus LUCIFER 1.5, and LUCIFER has the -1.5 deg LEFTIAA term which isn't nearly large enough to cause this effect.

We lost more than half the night to various telescope failures. One of the chillers in the utility bldg failed. We limped along through the second half without mirror ventilation and with big fans blowing hot air out of the computer room. We got some WFS checking done. GCS active optics does now collimate the telescope on/off-axis. MCSPU (Jet) crashed 3 times losing another 3 hours of time. There were various other minor problems mixed in with these. We had lots of problems finding the star. Partly this is due to the changing temperature gradient of the telescope structure, partly due to many restarts of MCSPU and PCS, and partly due to the problem with the off-axis probe position for LUCIFER.

Nighttime Activities

After many days of rain, the sky is clearing off.

00:05 Start mirror ventilation with manual setpoint of 9 degC.

01:08 Open facing east for pre-sunset sky flats. T=7degC, D=5degC, wind light from SW, sky mostly clear with dissipating cumulus clouds.

02:01 Close for fog / low clouds blowing over the mountain.

02:20 Fog is gone, and humidity is down, so we prepare to reopen as soon as Nancy and Peter have checked the window ventilation.

02:28 Open shutter doors again. T=7 degC D=-3 degC.

02:35 First preset fails to find star

02:42 Send again with typo in the guide star coordinates corrected.

02:49 Move guide probe on-axis and adjust IE=-107 CA=-46

02:52 Restart GCSL after John's still running readGuideCam loop crashed it.

02:53 Back to Nancy's original off-axis preset. A star is there on-axis.

03:00 Adjust IE/CA on the off-axis star (maybe not a good plan) IE=-118 CA=-42

03:05 Preset in guide mode, fails with jump?

03:07 Preset fails for OSS not running. Then still can't find the star that we see in the acquisition image.

03:11 Preset to guide on-axis. no star

03:12 Preset to guide on-axis again with IE=-107 CA=-46 - Conclude that our initial centering was on the wrong star.

Restore IE=-90 CA=-60

03:19 Find a new star on-axis IE=-123 CA=-83

03:20 Preset Active on-axis. WFS is saturated, setWFSExpTime 15000

03:25 Preset Active off-axis. no star

Problem with position probe on off-axis star

03:32 Preset back to the on-axis star, and let WFS converge. LUCIFER is taking images at each position.

03:35 Preset 30 arcsec South, using the same star for guiding. (found star at 351, 359)

03:40 Preset 60 arcsec South, but star is now up in the corner of the acquisition image. Vibration made some cigar images for some seconds.

One of the 45degF chillers has tripped off. David resets it several times during the evening.

03:50 Preset 90 arcsec South, star at 471,473, missed on first guide image, but pulled it in. Preset succeeded despite the following error.

Sep 12 20:40:54 lbtmu107 LBT_left_GCS: TCPSocket::recv error: Transport endpoint is not connected

03:57 Preset 120 arcsec South, star is just tucked in the corner of the acquisition image, but the guider picks it up.

04:05 LBTO@lbtmu107:24 % stopWFS WFS loop NOT stopped! but succeed on second try.

04:06 Doug running IDL WFS

04:09 Sent corrections to primary /Repository/AGW_Data/080913w, exp = 15 sec, wfsc 2, 3, 4,5

04:13 exp = 10 sec, 6, 7, 8 Looks like a bunch of astigmatism and coma. Try a BS star

04:14 Set mirror setpoints back to automatic (PID) mode

04:17 Authorize IRTC

04:25 pointto BS9178, TRACK, John search for star, found at offset -30 -30, and set IE=-40 and CA=-50. We think that the pointing offsets are changing rapidly because the telescope structure is out of temperature equilibrium.

04:40 wfsc 9-15

04:49 TO apply +1000 Z7: see -600 Z7 on wfsc #16

04:51 apply corrections wfsc 17-21 Still bad

04:55 Adjust focus in AIP_L.cfg from 21 to 29 (our previous setting was overwritten by the build)

05:03 Preset to a new star

05:12 My third IRTC.offset command failed because pause guiding didn't work. Why didn't the first two fail? PA on Guider looks OK.

05:17 pointto GUIDE, found star at 286,213

David reset 45degF chiller for the third time.

05:27 apply z15 -200 see ~+100 change on wfsc to ~0

05:30 apply z14 100 see +40 change to +80

05:34 apply z14 -300 (-200 from original) see

05:37 -300 doesn't look better

05:40 -300 doesn't look better, driving wrong axis, z14 -210

05:42 +500 z14 -60

05:45 +800 z14 +140

05:49 -500 z14 to 6

05:50 z25 +200 can not tell difference #37

05:55 z25 -400 can not tell difference #38

05:57 back to original wfsc 39-42

Took LUCIFER image here

06:00 BS9183 gstar=51, didn't find guide star

06:01 TRACK, found star at 455,128, John reset IE=-29 CA -45

06:04 didn't receive full CCD readout

06:09 found star off-axis at 174,215

Nancy confirms that LUCIFER position angle rotation is correct with LEFTIAA = -1.533.

LUCIFER image luci_20080912_0110 has 0.7 arcsec images.

06:xx on-axis ??

06:17 guiderimage004215.fits off-axis GUIDE, gstar=51, not subtracting off-axis corrections. Changed IDL routines.

06:20 off-axis gstar=51, with off-axis corrections applied. collimate 49-55, found star at 166,231

Note that guide images have an upper left to lower right trend. Elevation? Vibration?

06:28 go on-axis and measure aberrations 56-58

06:35 on-axis with "ACTIVE" /tmp/wfscimage0000509-521

06:46 off-axis with "ACTIVE" 522-534, found star at 138,253 at probe position x=

LUCIFER image luci_20080912_0115

06:55 on-axis with "GUIDE" wfsc 59-61 to check the GCS off-axis result

07:02 off-axis with "ACTIVE" 0000535-545, found star at 163,115 with probe position x=102 y=513.8

LUCI image 117 is out of focus.

Beware that we have accidentally been sending the active wavefront corrections as default. Apparently all the copies of PSFGUI don't syncronize the radio buttons.

07:12 Authorize LUCIFER

07:15 LUCI preset to off-axis star fails to find any star at probe position x=24.6 y=527.0

07:19 LUCI preset to on-axis star fails with full MCSPU shutdown - even our window disappeared. jet has rebooted on its own as of 00:26 MST.

07:30 David has to restart the jet rpcserver (as is common after a reboot)

07:38 Back to preset (lost 20 minutes)

07:42 Back to Nancy's original object in GUIDE mode, but acquire fails, manual adjust Z global offset by 0.1 to focus.

07:47 now guiding works and locks onto one of two stars in the acquisition image.

07:52 manually dialing Active Optics corrections out of M2 while GCS WFS loop puts them into M1.

07:58 Stop SX mirror ventilation to try to keep alive longer on only one chiller

08:01 stop guiding

08:03 resend the LUCI preset, but MCSPU crashes again. Lost only 12 minutes this time, since we had a more clear idea what happened, but then had trouble getting PCS to run correctly after restarting it so lost another 5 minutes.

Called Dan Cox (about jet)

Called Butch and John L. (about chiller),

08:16 Resend LUCIFER preset. From the sound of building moving, slew polynomials made a mistake. Failed to acquire the guide stars

08:24 Resend LUCIFER preset with PA180 - they are checking for mask alignment. (could not reach the same guide star)

08:30 Jet crashes again.

08:48 Norm presses reset button in Jet after the core dump to lbtmu102 was unsuccessful due to NMIs.

08:56 Snapshot of telescope temps (chamber was ~15 degC 10 hours ago)

Location Air Steel
C-Ring Upper 6.42 6.92
C-Ring Middle 7.53 7.97
C-Ring Lower 7.25 7.72
Extension Left 5.65 5.25 (reversed?)
Extension Right 5.43 5.52
WindBrace Upper 5.02 5.95
WindBrace MiddleLeft 5.19 6.18
WindBrace MiddleRight 5.37 6.19
WindBrace LowerRight 6.42 7.08
MirrorCell LowerRight 6.17 6.47

09:00 Summer chiller CH0401 is stable at LWT of -8.44 degC (normal), but A/C in control / computer room is still overwhelmed.

09:03 Jet back online - lost 33 minutes.

09:06 LUCI preset failed because OSS not running.

09:08 LUCI preset again, failed to find off-axis star

09:14 pointto BS9188 gs=0 PA=0 TRACK move probe manually

09:18 spiral search found at -30,60 (IE CA had reverted to pointing model when we restarted PCS)

09:32 MSCPU crashed at offset -12 30 with IE=-65 CA=-25 (offset command eventually timed out in 300s)

09:42 LUCIFER taking pinhole image while telescope down, to check internal aberrations. See LUCI image 0135. Pinhole spots look OK.

10:05 Jet comes briefly back to life, but crashes again before we could even move the telescope. David goes up to give it the full power cycle reboot.

Call Dave A. to learn howto reload rotator FPGA

Problems with frost on LUCI window

11:26 pointto BS9107 Binary IE=-65 CA=-40

11:33 pointto BS9107 ACTIVE (but forgot to stop the manual images)

Sep 13 04:33:23 lbtmu107 LBT_left_GCS: TCPSocket::recv error: Bad file descriptor

11:37 pointto again after killing GCSL

setExpTime 800

Sep 13 04:39:06 lbtmu107 LBT_left_GCS: fatal error: did not receive full CCD readout from guide camera within timeout period of 6459 ms.

11:41 pointto ACTIVE - finally succeeds, starting WFSing

LUCI image 0137

11:52 pointto again, guiding on faint star

12:00 Lucifer preset

12:04 offset -30 60 fails after the spiral.cl script was working.

12:11 IE=-35, CA=-40 from centering on LUCIFER

12:12 LUCI preset finds on-axis star 361,417

12:14 LUCI preset with guidestar 60 arcsec South

Sep 13 05:13:50 lbtmu107 LBT_left_GCS: fatal error: did not receive full CCD readout from guide camera within timeout period of 6159 ms.
happened again around 05:17

12:18 LUCI preset tracking

12:20 object moved north along columns on offset, but "E-W offsets are wierd". (confused by frost and probe)

12:29 trying N-S offsets again

12:32 trying E-W offsets again. LUCIFER offsets (from the guide paddle) work correctly in TRACK mode.

12:35 LUCI preset in guide mode, but sky too bright for this star

12:40 Close dome. T=5degC D=3degC

-- JohnHill - 13 Sep 2008
Topic revision: r7 - 14 Sep 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback